OpenStudy (anonymous):

Find the slope of a line that passes through (–2, –3), and (1, 1). (1 point) onethird 1 2 fourthirds need serious help

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

serious help?

OpenStudy (anonymous):

im usually great math but this i cant seem to get

OpenStudy (misty1212):

from -2 to 1 on the x axis is right 3 units from -3 to 1 on the y axis is up 4 right 3, up 4, slope is \(\frac{4}{3}\)

OpenStudy (misty1212):

or you can do it the math teacher way \[\frac{1-(-3)}{1-(-2)}=\frac{1+3}{1+2}=\frac{4}{3}\]

OpenStudy (anonymous):

do you mind helping me with a few more

OpenStudy (misty1212):

math teachers \(\heartsuit\) the point slope formula

OpenStudy (misty1212):

sure

OpenStudy (anonymous):

For the equation –4y = 8x, what is the constant of variation? (1 point) –4 –2 1 2

OpenStudy (misty1212):

i got the question marks

OpenStudy (misty1212):

ok \[-4y=8x\\ y=-2x\]

OpenStudy (anonymous):

sorry my stupid laptops doing it

OpenStudy (misty1212):

constant of variation is \(-2\)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Suppose y varies directly with x, and y = 24 when x = 8. What is the value of y when x = 10? (1 point) 10 30 80 240